# GeoZero: Incentivizing Reasoning from Scratch on Geospatial Scenes

### v1

@@ 4 Experiment
In this section, we conduct comprehensive quantitative and qualitative experiments to thoroughly evaluate the capabilities of GeoZero. We first present its performance across four typical remote sensing vision–language tasks: scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC), as well as under challenging benchmarks. We then perform a series of ablation studies to validate the effectiveness of the constructed datasets and the proposed RL algorithm. Furthermore, we explore the relationship between the emergence of the thinking trajectory and model performance. Finally, we visualize several examples of reasoning texts generated by GeoZero, demonstrating its unique ability to perform emergent reasoning on geospatial scenes.
@@
Method UCM [ucm] AID [aid] RESISC-45 [nwpu] EuroSAT [eurosat] SIRI-WHU [siri-whu] WHU-RS19 [whurs19]
Qwen2.5-VL-7B-Instruct [qwen25-vl] 68.81 64.45 64.32 29.67 60.00 79.61
Qwen3-VL-8B-Instruct [qwen25-vl] 75.71 71.40 67.63 47.43 61.46 86.89
GeoChat [kuckreja2024geochat] 84.43 72.03 89.37 - 53.04 85.30
GeoMag [geomag] - 83.03 77.62 - - -
LHRSBot-Nova [li2025lhrs] - 88.32 86.80 63.54 72.21 95.63
VHM [pang2025vhm] - 91.70 94.54 - 70.88 95.80
SkySenseGPT [zhan2025skyeyegpt] - 92.25 - - 68.38 97.02
Earthdial [soni2024earthdial] 92.42 88.76 - - - 96.21
ScoreRS [scorers-r1] - 85.90 91.59 - 74.75 96.30
TinyRS-R1 [koksal2025tinyrs] - 90.20 92.90 - 76.80 95.60
RingMo-Agent [ringmoagent] 88.00 91.67 94.72 - - 96.81
Earth-Agent [earthagent] - 93.42 - - - 93.19
GeoZero (w/o RFT) 93.81 92.55 87.41 85.57 74.38 96.12
GeoZero (w/ RFT) 95.48 97.30 93.54 91.57 86.67 98.06
@@
Method RSVG [rsvg] DIOR-RSVG [dior-rsvg] VRS-VG [li2024vrsbench]
InternVL-2.5-8B-Instruct [internvl-25] - 15.39 -
Qwen2.5-VL-7B-Instruct [qwen25-vl] 11.41 34.32 39.31
Qwen3-VL-8B-Instruct [qwen25-vl] 5.95 3.50 0.08
GeoChat [kuckreja2024geochat] 14.67 19.77 -
SkySenseGPT [zhan2025skyeyegpt] - 12.87 -
LHRSBot-Nova [li2025lhrs] - 31.51 -
VHM [pang2025vhm] - 56.17 -
ScoreRS-R1 [scorers-r1] - 64.52 -
Earth-Agent [earthagent] - 60.46 -
TinyRS-R1 [koksal2025tinyrs] - 74.90 -
GeoGround [zhou2024geoground] 26.65 77.73 66.04
GeoZero (w/o RFT) 37.16 75.67 67.97
GeoZero (w/ RFT) 50.04 79.43 73.32
4.1 Experimental Settings
We train GeoZero following a two-stage SFT–RL framework, with Qwen3-VL-8B-Instruct [qwen25-vl] serving as the base model. Both SFT and RL are trained for 1 epoch. By controlling the gradient accumulation steps, we set the global batch size to 64 and 48, respectively. The learning rates for SFT and RL are 1e-4 and 5e-6, respectively, with weight decays of 0.1 and 0.01. During both stages, the vision encoder is kept frozen, while the remaining modules are fine-tuned using LoRA [lora] with a rank of 16 and $\alpha=32$. We employ bfloat16 precision and DeepSpeed [deepspeed] to reduce GPU memory consumption, and use FlashAttention [dao2022flashattention] to accelerate computation. For A2GRPO, the group number $G$ is set to 8, the temperature is 0.9, and $\beta$ in Equation (4) is 0.04. For AMTR, the gating parameters $k$ and $\tau$ are 12 and 0.5. The coefficient $\lambda$ in Equation (3) is set to 0.3. All experiments are conducted on 8 NVIDIA A100 GPUs.
4.2 Comparison to State-of-the-art MLLMs
Before presenting the results, we note that GeoZero-Raw contains training samples from several public datasets, while evaluations are performed on their respective test sets to ensure fairness and to assess generalization on unseen data. To further enhance reasoning in specific scenarios, we also apply reinforcement fine-tuning (RFT) with A2GRPO on individual datasets, following the practices in [visual-rft]. The detailed hyperparameter configurations of RFT for each dataset are provided in the supplementary material.
@@ 4.2 Comparison to State-of-the-art MLLMs
Method RSVQA-HR-Test1 RSVQA-HR-Test2
Presence Compare Presence Compare
InternVL-2.5-8B-Instruct [internvl-25] - - 65.80 75.50
Qwen2.5-VL-7B-Instruct [qwen25-vl] 64.04 75.61 62.11 74.24
Qwen3-VL-8B-Instruct [qwen25-vl] 65.58 75.79 66.04 75.39
GeoChat [kuckreja2024geochat] 56.36 82.74 58.45 83.19
SkyEyeGPT [zhan2025skyeyegpt] 84.95 85.63 83.50 80.28
GeoMag [geomag] 87.04 86.81 77.62 88.94
VHM [pang2025vhm] - - 64.00 83.50
SkySenseGPT [zhan2025skyeyegpt] - - 69.14 84.14
LHRSBot-Nova [li2025lhrs] - - 87.60 89.30
Earthdial [soni2024earthdial] - - 58.89 83.11
RingMo-Agent [ringmoagent] - - 75.24 83.92
TinyRS-R1 [koksal2025tinyrs] - - 68.60 73.50
ScoreRS [scorers-r1] - - 75.80 87.30
GeoZero (w/o RFT) 75.62 84.71 74.46 83.59
GeoZero (w/ RFT) 91.16 91.01 88.00 89.83
Typical Remote Sensing Vision–Language Tasks. Tables 25 present the performance comparison between GeoZero and existing state-of-the-art methods on the SC, VG, VQA, and IC tasks. For evaluation, the performance of SC and VQA is measured by accuracy, while VG adopts Acc@0.5, where a prediction is considered correct if the IoU between the predicted box and the ground truth exceeds 0.5. For IC, multiple metrics are employed for comprehensive evaluation, including BLEU-4, CIDEr, METEOR, ROUGE-1, and ROUGE-L.
@@
Method BLEU-4 CIDEr METEOR ROUGE-1 ROUGE-L
UCM-Captions Dataset [ucm-captions]
GeoChat [kuckreja2024geochat] - 14.27 14.40 13.22
SkyEyeGPT [zhan2025skyeyegpt] 78.41 236.75 46.24 - 79.49
Earthdial [soni2024earthdial] - - 51.42 40.00 34.15
RingMo-Agent [ringmoagent] 77.63 373.68 51.79 - 85.51
GeoZero (w/o RFT) 55.80 279.11 76.52 76.30 72.77
GeoZero (w/ RFT) 75.72 393.57 83.75 84.60 83.75
Sydney-Captions Dataset [ucm-captions]
GeoChat [kuckreja2024geochat] - - 10.63 12.00 11.26
SkyEyeGPT [zhan2025skyeyegpt] 77.40 181.06 46.62 - 77.74
Earthdial [soni2024earthdial] - - 57.31 49.39 41.00
GeoZero (w/o RFT) 34.38 125.12 51.72 54.36 48.54
GeoZero (w/ RFT) 64.08 293.49 75.12 75.72 74.12
NWPU-Captions Dataset [nwpu-captions]
GeoChat [kuckreja2024geochat] - - 15.21 14.86 12.54
Earthdial [soni2024earthdial] - - 80.61 45.84 39.96
RingMo-Agent [ringmoagent] - 48.39 18.81 - 34.11
GeoZero (w/o RFT) 54.09 167.97 72.75 75.28 69.69
GeoZero (w/ RFT) 40.26 142.98 65.53 67.98 63.10
RSICD Dataset [rsicd]
GeoChat [kuckreja2024geochat] - 12.39 13.48 11.59 -
SkyEyeGPT [zhan2025skyeyegpt] 59.99 83.65 35.35 - 62.63
Earthdial [soni2024earthdial] - - 56.18 33.77 27.61
GeoZero (w/o RFT) 29.17 97.45 48.18 55.94 47.33
GeoZero (w/ RFT) 35.35 114.82 52.31 60.16 54.47
For the SC task, following the evaluation protocol in [pang2025vhm, li2025lhrs, koksal2025tinyrs, scorers-r1], we adopt six widely used datasets for evaluation. As shown in Table 2, our method surpasses recent agent-based models [ringmoagent, earthagent] on most benchmarks, achieving state-of-the-art performance. Notably, even without RFT, GeoZero still exhibits highly competitive results, particularly on the unseen UCM dataset [ucm]. For the VG task, we employ RSVG [rsvg], DIOR-RSVG [dior-rsvg], and VRS-VG [li2024vrsbench] datasets. We observe that GeoZero demonstrates clear superiority, significantly outperforming previous remote sensing MLLMs, including the specialized grounding model GeoGround [zhou2024geoground]. For the VQA task, we adopt the classical RSVQA-HR dataset [rsvqa], which contains two test sets. The results in Table 4 show that our model achieves the best accuracy among all methods, while maintaining comparable performance even without RFT. Finally, for the IC task, we evaluate GeoZero on four datasets, including UCM-Captions [ucm-captions], Sydney-Captions [ucm-captions], NWPU-Captions [nwpu-captions], and RSICD [rsicd]. As shown in Table 5, our model consistently achieves the best results across multiple evaluation metrics, particularly in terms of the CIDEr, where GeoZero outperforms all comparison approaches despite CIDEr not being explicitly optimized within the answer reward.
Challenging Benchmarks. Beyond these classical tasks, we further evaluate GeoZero on two recently proposed challenging benchmarks: XLRS-Bench [Wang_2025_CVPR_XLRSBench] and CHOICE [an2025choice]. Specifically, XLRS-Bench features ultra–high-resolution imagery with rich semantic contexts across large-scale scenes, while CHOICE comprises multi-source data covering multiple modalities and platforms with global geographic distribution. More importantly, both benchmarks include multiple subtask evaluations across perception and reasoning dimensions, providing a comprehensive assessment of GeoZero’s intrinsic cognition toward geospatial scenes. As shown in Table 6, even under complex and challenging conditions, our model achieves consistently strong performance compared with existing remote sensing and general-purpose MLLMs. This improvement can be attributed to the enhanced reasoning capability acquired during the RL stage, where training on hard samples enables better adaptation to difficult benchmarks.
Table 6: Performance comparison of different MLLMs on comprehensive benchmarks. The average accuracy of L3 subtasks is used as the metric. The best and second-best results are highlighted in bold and blue, respectively. *: Training with HighRS dataset [wang2025geollavak], where only images smaller than 1024×1024 are used due to GPU memory constraints.
Method XLRSBench [Wang_2025_CVPR_XLRSBench] CHOICE [an2025choice]
GeoChat [kuckreja2024geochat] 22.90 52.80
GPT-4o [hurst2024gpt4o] 32.40 68.02
GeoZero (w/o RFT) 47.24 71.67
GeoZero (w/ RFT)* 48.10 72.06
@@ 4.3 Ablation Study
To validate the effectiveness of the constructed datasets and the proposed RL algorithm, we select two representative tasks, SC and VG, using the UCM and RSVG datasets, respectively.
Training Paradigm. We first examine the impact of different training strategies, as summarized in Table 7. It can be observed that direct RL training (RL only) can indeed trigger reasoning generation ($\bullet$). However, due to the lack of fundamental geospatial understanding without SFT training, the model suffers from severe hallucinations and extremely poor accuracy, even after subsequent RFT. In contrast, performing only SFT or SFT+RFT yields strong accuracy but fails to activate reasoning ($\circ$). This can be attributed to the absence of CoT supervision in SFT and the limited scale and diversity of downstream datasets (compared with GeoZero-Hard) in RFT, which provide insufficient signals for reasoning to emerge. By incorporating the constructed hard-sample pool, GeoZero-Hard, into the SFT–RL training paradigm (GeoZero w/o RFT), the model not only maintains high accuracy but also successfully activates reasoning, demonstrating its ability to think from scratch. Furthermore, after applying RFT on downstream datasets (GeoZero w/ RFT), both reasoning consistency and task accuracy are further improved, demonstrating the effectiveness of our training paradigm in incentivizing reasoning and enhancing geospatial understanding.
@@
Qwen3-VL-8B-Instruct 75.71  ($\circ$) 5.95  ($\circ$)
@@
GeoZero (w/o RFT) 93.81  ($\bullet$) 37.16  ($\bullet$)
GeoZero (w/ RFT) 95.48  ($\bullet$) 50.04  ($\bullet$)
Table 8: Ablation study on the dataset used during RL. “Random” means the data are randomly selected from GeoZero-Raw, while the construction method of the SFT dataset is the same as GeoZero-Instruct. $\bullet$ and $\circ$ follow the same definitions as in Table 7.
@@
GeoZero (w/o RFT) Random 94.05  ($\circ$) 18.34  ($\circ$)
GeoZero (w/o RFT) GeoZero-Hard 93.81  ($\bullet$) 37.16  ($\bullet$)
GeoZero (w/ RFT) Random 95.24  ($\circ$) 46.70  ($\bullet$)
GeoZero (w/ RFT) GeoZero-Hard 95.48  ($\bullet$) 50.04  ($\bullet$)
@@
GeoZero (w/o RFT) 91.67  ($\circ$) 26.49  ($\circ$)
GeoZero (w/o RFT) 93.10  ($\circ$) 28.28  ($\circ$)
GeoZero (w/o RFT) 93.57  ($\circ$) 26.57  ($\circ$)
GeoZero (w/o RFT) 93.81  ($\bullet$) 37.16  ($\bullet$)
GeoZero (w/ RFT) 93.81  ($\circ$) 33.01  ($\circ$)
GeoZero (w/ RFT) 95.24  ($\circ$) 49.80  ($\circ$)
GeoZero (w/ RFT) 94.76  ($\bullet$) 47.03  ($\bullet$)
GeoZero (w/ RFT) 95.48  ($\bullet$) 50.04  ($\bullet$)
@@ 4.4 Visualization
Figure 4: Visualization of reasoning processes generated by GeoZero across different tasks. Examples are drawn from the test sets of AID, DIOR-RSVG, and RSVQA-HR. Red text highlights the model’s reasoning steps leading to the final answers. GT: ground truth.
Thinking v.s. Answering. To further investigate how reasoning influences final predictions, we analyze the relationship between answer accuracy and both reasoning length and reasoning quality score $s_{t}$. Specifically, all samples are grouped based on their reasoning length and quality score, and the average accuracy within each group is computed. As shown in Fig. 3(a), accuracy increases with longer reasoning chains, indicating that extended reasoning generally benefits answer correctness. However, when the reasoning becomes excessively long, i.e., overthinking, the performance slightly drops. Similarly, Fig. 3(b) reveals a positive correlation between $s_{t}$ and answer accuracy, implying that higher-quality reasoning contributes to more reliable decisions. These results demonstrate that GeoZero’s reasoning process plays a causal role in guiding accurate answers, validating the effectiveness of our proposed method.
Qualitative Reasoning Results. Finally, to gain an intuitive understanding of the model’s reasoning process, we qualitatively visualize several examples of GeoZero’s predictions across different tasks, as shown in Figure 4. It can be observed that before producing the final answer, GeoZero explicitly articulates its internal reasoning by describing key visual attributes such as object shape, color, spatial layout, quantity, and contextual relationships. These examples demonstrate that GeoZero is capable of generating coherent and interpretable reasoning trajectories, showcasing its ability to perform emergent reasoning over complex geospatial scenes, while also highlighting its potential for building transparent and explainable AI systems.

### v2

@@ 4 Experiment
In this section, we evaluate GeoZero through both quantitative and qualitative experiments. We first report results on four typical remote sensing vision-language tasks: scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC). We then conduct ablation studies to analyze the contribution of our datasets and RL algorithm, followed by an analysis of the correlation between thinking behaviors and answer correctness. Finally, we visualize representative reasoning examples to illustrate GeoZero’s emergent geospatial reasoning.
4.1 Experimental Settings
We train GeoZero following a two-stage SFT–RL framework, with Qwen3-VL-8B-Instruct [4] serving as the base model. Both SFT and RL are trained for 1 epoch. By controlling the gradient accumulation steps, we set the global batch size to 64 and 48, respectively. The learning rates for SFT and RL are 1e-4 and 5e-6, respectively, with weight decays of 0.1 and 0.01. During both stages, the vision encoder is kept frozen, while the remaining modules are fine-tuned using LoRA [20] with a rank of 16 and $\alpha=32$. We employ bfloat16 precision and DeepSpeed [49] to reduce GPU memory consumption, and use FlashAttention [12] to accelerate computation. For A2GRPO, the group number $G$ is set to 8, the temperature is 0.9, and $\beta$ in Equation (4) is 0.04. For AMTR, the gating parameters $k$ and $\tau$ are 12 and 0.5. The coefficient $\lambda$ in Equation (3) is set to 0.3. All experiments are conducted on 8 NVIDIA A100 GPUs.
4.2 Comparison to State-of-the-art MLLMs
@@
Method UCM AID
Qwen3-VL-8B-Instruct [4] 75.71 71.40
GeoChat [28] 84.43 72.03
VHM [47] - 91.70
ScoreRS [45] - 85.90
RingMo-Agent [21] 88.00 91.67
TinyRS-R1 [27] - 90.20
GeoZero 93.81 92.55
@@
Method RSVG DIOR-RSVG
Qwen3-VL-8B-Instruct [4] 26.41 51.51
GeoChat [28] 14.67 19.77
VHM [47] - 56.17
ScoreRS-R1 [45] - 64.52
TinyRS-R1 [27] - 74.90
GeoGround [74] 26.65 77.73
GeoZero 37.16 75.67
@@
Method RSVQA-HR
Presence Compare
Qwen3-VL-8B-Instruct [4] 66.04 75.39
GeoChat [28] 58.45 83.19
VHM [47] 64.00 83.50
TinyRS-R1 [27] 68.60 73.50
GeoZero 74.46 83.59
@@
Method RSICD
BLEU-4 CIDEr METEOR
GeoChat [28] - 12.39 13.48
SkyEyeGPT [68] 59.99 83.65 35.35
GeoZero 29.17 97.45 48.18
Typical Remote Sensing Vision-Language Tasks. Tables 2-5 summarize the performance of GeoZero against representative state-of-the-art remote sensing MLLMs on four typical tasks: SC, VG, VQA, and IC. For evaluation, SC and VQA are measured by accuracy. VG adopts Acc@0.5, where a prediction is considered correct if the IoU between the predicted box and the ground truth exceeds 0.5. For IC, we report standard captioning metrics, including BLEU-4, CIDEr, and METEOR.
For SC, we evaluate on two widely used benchmarks, UCM [65] and AID [61], following common evaluation practices in recent works [47, 27, 45]. As shown in Table 2, GeoZero achieves superior performance on both datasets, outperforming both general MLLMs and recent remote sensing MLLMs. Notably, GeoZero yields a clear margin on UCM, indicating strong generalization to unseen scenes.
For VG, we report results on RSVG [55] and DIOR-RSVG [67]. As shown in Table 3, GeoZero delivers the best performance on RSVG and remains highly competitive on DIOR-RSVG, approaching the specialized grounding model GeoGround [74]. These results demonstrate that our method effectively strengthens spatial localization and grounding capability.
For VQA, we adopt RSVQA-HR [39] for evaluation and report accuracy on its two question types (Presence and Compare). The results in Table 4 show that GeoZero attains the highest accuracy on both question types, improving substantially over the general-purpose base model while remaining competitive against prior remote sensing MLLMs.
Finally, for IC, we evaluate on RSICD [40]. As shown in Table 5, GeoZero achieves the best CIDEr and METEOR among the compared methods, despite CIDEr not being explicitly optimized in the task-specific answer reward. This suggests that the proposed training framework can improve the overall quality of generated descriptions beyond the directly optimized components.
@@ 4.3 Ablation Study
To validate the effectiveness of the constructed datasets and the proposed RL algorithm, we conduct ablation studies on two representative tasks: SC and VG, using the UCM and RSVG datasets, respectively.
Notably, to further analyze the activation of reasoning in task-specific scenarios, in addition to the standard SFT and RL stages, we additionally introduce reinforcement fine-tuning (RFT) with A2GRPO on each dataset (denoted as GeoZero+RFT). Specifically, RFT is performed only on the training split of each dataset, with no overlap with the evaluation data. The detailed configurations of RFT are provided in Section H of the supplementary material.
@@
Qwen3-VL-8B-Instruct 75.71  ($\circ$) 26.41  ($\circ$)
@@
GeoZero 93.81  ($\bullet$) 37.16  ($\bullet$)
GeoZero+RFT 95.48  ($\bullet$) 50.04  ($\bullet$)
Table 7: Ablation study on the dataset used during RL. “Random” means the data are randomly selected from GeoZero-Raw-val (see Section-D in the supplementary material), while the construction method of the SFT dataset is the same as GeoZero-Instruct. $\bullet$ and $\circ$ follow the same definitions as in Table 6.
@@
GeoZero Random 94.05  ($\circ$) 18.34  ($\circ$)
GeoZero GeoZero-Hard 93.81  ($\bullet$) 37.16  ($\bullet$)
GeoZero+RFT Random 95.24  ($\circ$) 46.70  ($\bullet$)
GeoZero+RFT GeoZero-Hard 95.48  ($\bullet$) 50.04  ($\bullet$)
@@
GeoZero 91.67  ($\circ$) 26.49  ($\circ$)
GeoZero 93.10  ($\circ$) 28.28  ($\circ$)
GeoZero 93.57  ($\circ$) 26.57  ($\circ$)
GeoZero 93.81  ($\bullet$) 37.16  ($\bullet$)
GeoZero+RFT 93.81  ($\circ$) 33.01  ($\circ$)
GeoZero+RFT 95.24  ($\circ$) 49.80  ($\circ$)
GeoZero+RFT 94.76  ($\bullet$) 47.03  ($\bullet$)
GeoZero+RFT 95.48  ($\bullet$) 50.04  ($\bullet$)
Training Paradigm. We first examine the impact of different training strategies, as summarized in Table 6. Direct RL training (RL only) can trigger reasoning generation ($\bullet$). However, without SFT, the model lacks fundamental geospatial knowledge and performs poorly on challenging benchmarks (e.g., RSVG). Although RFT can partially recover accuracy, it still falls short of the SFT-RL pipeline trained with GeoZero-Hard. In contrast, SFT-only and SFT+RFT achieve strong accuracy but fail to activate reasoning ($\circ$), likely due to the absence of CoT supervision in SFT and the limited scale and diversity of downstream training data (compared with GeoZero-Hard) in RFT, which provide insufficient optimization signals to induce reasoning. By incorporating the constructed hard-sample pool, GeoZero-Hard, into the SFT-RL paradigm (GeoZero w/o RFT), the model maintains high accuracy while consistently activating reasoning, indicating its ability to think from scratch. Further applying dataset-specific RFT (GeoZero w/ RFT) improves both task accuracy and reasoning consistency, validating the effectiveness of our training paradigm.
@@ 4.4 Visualization
Thinking v.s. Answering. To further investigate how reasoning relates to final predictions, we analyze the relationship between answer accuracy and both reasoning length and reasoning quality score $s_{t}$. Specifically, all samples are grouped based on their reasoning length and quality score, and the average accuracy within each group is computed. As shown in Figure 3(a), accuracy increases with longer reasoning chains, indicating that extended reasoning generally benefits answer correctness. However, when the reasoning becomes excessively long, i.e., overthinking, the performance slightly drops. Similarly, Figure 3(b) reveals a positive correlation between $s_{t}$ and answer accuracy, indicating that samples with higher reasoning quality scores tend to achieve higher accuracy. Overall, these results suggest a strong association between reasoning quality and task performance, supporting our design choice of encouraging explicit reasoning during RL. More fine-grained analyses are provided in Section-I of the supplementary material.
Figure 4: Visualization of reasoning processes generated by GeoZero on different tasks. Examples are drawn from the test sets of AID and DIOR-RSVG. Red text highlights the model’s reasoning steps leading to the final answers. GT: ground truth.
Qualitative Reasoning Results. Finally, to gain an intuitive understanding of the model’s reasoning process, we qualitatively visualize several examples of GeoZero’s predictions, as shown in Figure 4. It can be observed that before producing the final answer, GeoZero explicitly articulates its internal reasoning by describing key visual attributes such as object shape, color, spatial layout, quantity, and contextual relationships. These examples demonstrate that GeoZero is capable of generating coherent and interpretable reasoning trajectories, showcasing its ability to perform emergent reasoning over complex geospatial scenes, while also highlighting its potential for building transparent and explainable AI systems. Additional reasoning examples are shown in Section-J of the supplementary material.